The equation of a transverse wave traveling on a string is given by y A sin(kx - cot). Data: A-22 mm, k-13 rad/m, - 240 rad/s. What is the amplitude? Submit Answer Tries 0/99 What is the frequency? Submit Answer Tries 0/99 What is the wave velocity? Submit Answer Tries 0/99 What is the wavelen Submit Answer Tries 0/99 For the same wave, find the maximum transverse speed of a particle in the string. Submit Answer Tries 0/99
Answer #1

1) amplitude A =22 mm = 22*10-3 m

2) frequency f = \omega /2\pi = 240/(2\pi) = 38.19 Hz

3) velocity = \omega /k = 240/13 = 18.46 m/s

4) wavelength \lambda = 2\pi/k = 2\pi/13 = 0.483 m

5) For a particle undergoing simple hormonic motion , its max speed v = \omega A = 240 * 22*10-3 = 5.28 m/s